Consider an economy with three domestic agents: Ann, Bill and Carl. And a foreign economy with a representative agent Fer

.Ann owns land and uses the trees in their yard for wood. She cuts down the trees and sells the wood .Bill owns a firm that buys wood and build and sell tables. . Carl is a retired agent that has some wealth. There is no government. . During the year the following transactions happen: Ann sold 700 pounds of wood at price $10/pound. The other 300 pounds are kept as inventory of the firm. Bill bought 500 pounds of wood, produced and sold 100 tables at price $500/table. Fer bought 30 tables from Bill and 200 pounds of wood from Ann. The domestic agents (A, B & C) bought 70 tables from the firm owned by Bill Bills capital depreciated by $1000. o o o o o Calculate GDP for this economy and do the income decomposition (Wages, Taxes, Net Profits, Depreciation)
Answer #1

Solution:

Firm A produceS1000 pounds of wood (As intermediate goods sells 500 pounds sell to B, As final goods sell 200 pounds to others, 300 pounds sell to self as inventory) at $10 per pound.

Firm B produced 100 tables with the usage of 500 pounds of wood as inputs, and capital depreciated by $1,000). (Sold 70 tables to ABC as domestic consumption, Sold to F 30 tables as export) at $500 per table

 

The three approaches to calculate GDP are as follows:

1) Product approach (market value, newly produced, final goods and services)

GDP = (1,000 - 500) * $10 + (100 * $500) = $55,000

 

2) Expenditure approach:

GDP = C + I + G + NX;

C = (200*$10) + (70*$500) = $37,000

I = 300 * $10 = $3,000

G= $0

NX = 30*$500 - $0 = $15,000

Thus, GDP = $37,000 + $3,000 + $0 - $15,000 = $55,000

 

3) Income approach:

A's profit: 1,000 * $10 = $10,000

B's Net Profit: (100 * $500) - (500 * $10) - 1,000 = $44,000

Depreciated capital = $1,000

GDP = $10,000 + $45,000 + $1,000 = $55,000